OpenStudy (anonymous):

Suppose the carnival's organizer wants to reduce the chances players may win. He decides to reduce the area of the circular disk by half. What is the diameter of the new disk? The original diameter is 8 in.

OpenStudy (asadkarim7):

area of 1st disk = pi 4^2 = 50.3in ^2 50.3/2 = 25.13 in ^2 r^2= 25.13/pi r^2 = 8 r =2.8 d= 5.6 in